OpenStudy (anonymous):

x^2-9/x^2-49 divided by a+3/ a+7 i factored and got (x+3)(x-3)(a+3) all over (x+7)(x-7)(a+7) but now im stuck

OpenStudy (anonymous):

Well, u would be since 3's and 7's don't mix and neither do a's and x's. (By the way, the (a+7) should be on top)

OpenStudy (anonymous):

i dont understand, i factored them out can u help me?

OpenStudy (anonymous):

To divide by a fraction, turn it upside down and multiply by it instead.

OpenStudy (anonymous):

so one of the 7's go on top?

OpenStudy (anonymous):

Just do what I said.

OpenStudy (anonymous):

i will, i'm just confused about how to fully set it up

OpenStudy (anonymous):

I repeat, to divide by a fraction, turn the fraction upside down and multiply by it instead.

OpenStudy (anonymous):

take what u r dividing by. whatever is on the bottom put it on top. whatever was on top put it on the bottom. now multiply that in

OpenStudy (anonymous):

so its a+3/a+7?

OpenStudy (anonymous):

Yes!! U multiply by that...

OpenStudy (anonymous):

is a+3/a+7 the answer because they are in a fraction a+3 is over a+7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What r u talking about?

OpenStudy (anonymous):

Look, the answer you had in your question is right except that u have the (a+3) on top and the (a+7) on the bottom when it should be the other way round. Which it will be when you do what I said. What's 1 divided by 1/2 (a half)? Answer 2 (2 halves make 1). How to get 2? Turn 172 the other way up, 2/1. And multiply. 1*2/1 = 2. Try it with a few simple number fractions and you will soon see how it works. And it is the same thing for algebra.

OpenStudy (anonymous):

\[(A / B) / (C / D) = A / B * D/C = AD/BC\]